OpenStudy (anonymous):

Antonio is seven less than four times his daughter's age. Antonio is 37; how old is his daughter?

OpenStudy (anonymous):

let daughter's age = x years four times daughter's age = 4x then Antonio's age = (4x - 7) years ATQ, 4x - 7 = 37 => 4x = 37 + 7 => x = 44/4 => x = 11 So, his daughter is 11 years old (as we took her age as x)